# How Far Are We From AGI?

### v1

@@ 6.2 AGI Evaluation
The concept of evaluating AGI traces back to the famous Turing Test proposed by Alan Turing in 1950 (Turing, 1950). Turing posited that a machine could be considered intelligent if it could converse with a human so that the human could not distinguish whether they were conversing with a machine or another human. This laid the groundwork for the field of AGI evaluation.
@@ 6.2.1 Expectations for AGI Evaluation
It is undeniably challenging to design evaluation frameworks that satisfy all the recommended characteristics, but nonetheless, a well-constructed evaluation pipeline can help reflect the true power of increasingly sophisticated AGI systems.
@@
The new generation of model evaluations should focus on assessing AI systems across multiple dimensions, considering not only their ability to engage in human-like conversation but also their capacity to reason, learn, adapt, and solve complex problems. These tests should encompass a broader range of cognitive abilities and evaluate the AI’s performance in real-world scenarios. Drawing from the concepts of AGI internal, external, and system levels as we discussed in earlier sections, we can outline the key aspects that a new AGI evaluation framework should address:
@@
  • System level. Examine the AI’s overall behavior and its ability to pursue long-term goals, collaborate with humans and other AI systems, and make ethical decisions. This could involve scenarios that assess the AI’s alignment with human values, its transparency and explainability, and its robustness and reliability in uncertain and adversarial situations.
@@
As we will discuss in the next section, the current evaluation frameworks are far away from achieving what we expect (Team, 2023). Here we list a couple of challenges associated with AGI evaluation design, and we provide some concrete examples in each category:
@@
  • Long feedback loops. The trend toward more general-purpose AI indicates that models would gradually become more in people’s lives, making them more interactive. Instead of single-task solving, AGI systems will get multi-step feedback, often extending a longer period, making evaluation more complex. One commonly seen case happens in search engine development where a mature system needs to improve the click-through rate and track down users’ following actions such as purchases, comments, and after-sale satisfaction. Another potential situation appears when a health AI robot monitors a patient’s biological status constantly. Both of these require evaluations that span an elongated period of time.
  • Complicated environment setup. Many tasks inevitably presuppose more complicated environment setups such as robotic manipulation, self-driving assistance, and program synthesis. Scaling the evaluation of these applications necessitates a higher level of automatic environment generation, verification, and measurement. Often, the challenge associated with these tasks also comes with the difficulty of applying a single metric for comparison and can usually encounter many physical constraints that are hard to overcome.
  • Super-evaluation. Similar to super alignment, when AI models start to surpass humans, the evaluation becomes prohibitively more challenging. Taking the example of theorem proving, one can imagine that the real meaning of these models comes when they can prove unsolved theorems, which require greater expertise to verify. With the current setup, the framework might determine whether an AI agent can beat humans for a specific task (e.g., AlphaGo Silver et al. (2017) and  (Lake and Baroni, 2023)) but might not be as confident to access its absolute ability beyond.
@@ 6.2.2 Current Evaluations and Their Limitations
In this section, we summarize several representative works focusing on existing AI evaluation benchmarks. One category aims to provide a single pipeline suite consisting of many different tests such as AGIEVAL (Zhong et al., 2023a) and Huggingface Open LLM Leaderboard  [^9] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard for language understanding, reasoning, knowledge, and interaction abilities, and MT-BENCH (Zheng et al., 2024) for multi-task generalization ability. The GAIA benchmark (Mialon et al., 2023) constitutes a significant stride in this direction. It is specifically designed to evaluate General AI Assistants, presenting a series of real-world questions that test fundamental competencies such as reasoning, multi-modality handling, web browsing, and tool-use proficiency. AGENTBENCH is another comprehensive benchmark (Liu et al., 2024c) suite designed for evaluating the efficacy of LLMs as autonomous decision-making agents across eight interactive environments, highlighting the performance discrepancy between leading commercial models and open-source counterparts. On a granular level, there are many prior efforts in accessing a model’s specialized ability, which can be roughly divided into several aspects: accuracy, calibration and uncertainty, robustness, fairness, bias and stereotypes, toxicity, and efficiency (Liang et al., 2023), which can be further classified into two sets based on their objectives. OpenAGI (Ge et al., 2023b) is an open-source platform designed to advance AGI by integrating LLMs with domain-specific expert models. It utilizes a dual strategy of benchmark and open-ended tasks to evaluate.
@@
  • Creativity. Ever since the born of generative models, research has been ongoing into the use of them to model human creative processes, to mimic or complement them, in art, music, and literature (Cardoso et al., 2009; Colton et al., 2012). The creativity is mainly linked to the diversity of generated content with specific tasks like storytelling (Alhussain and Azmi, 2021) emerges. Recent works focus on understanding and prompting LLMs to generate creative textual content (Zhao et al., 2024c; Shanahan and Clarke, 2023).
@@
  • Lack of failure analysis. Almost all benchmarks we have seen so far give aggregated results, usually in the form of averages and percentiles. However, benchmarks should in principle provide more insights into improving a certain system. The most informative feedback would contain information about the analysis of the failing or worst cases. This can also showcase the potential pitfalls and risks associated with a specific system also help with the decision for danger-critical applications.
@@ 6.3 How to Get to the Next AGI Level
  • Advanced code machine brings the ultimate AGI within easy reach. We believe that if advanced code machine is applied to human society, the realization of the ultimate AGI is just around the corner. The AGI in level 1 and 2 is limited since they require a variety of data collected by humans and require specific optimization goals designed by humans when they tackle different tasks. Human-in-the-loop makes it impossible for AGI at these two levels to realize self-evolve. Advanced code machines solve the above challenges in two ways: 1) They enable AGI to interact with the real world and obtain large amounts of domain data. In the scenario of a single agent, an advanced code machine takes writing code as the most basic tool for AGI to interact with the world, enabling AGI to plan and reason in the form of code and get feedback on the real world through the results of code. When it comes to multi-agent scenarios, each agent can be regarded as a unique code machine based on their profile. Then, through the interactions between agents and the interaction between agents and the real world, AGI can obtain enough data for self-training and evolution. 2) It makes it possible for AGI to automatically define optimization goals for different tasks. With the ability to write codes, AGI can do try-and-error in the different tasks and obtain feedback through the interaction between code and the real-world environment. These feedbacks contain information about how well the AGI has completed its tasks and can take many forms, which can e differentiable or non-differentiable. Some techniques based on reinforcement learning can be introduced to use these different forms of feedback to align and self-evolve the AGI. In this case, the evolution of stronger AGI can be achieved without requiring human to specify goals.
@@
  • Super realistic simulation promotes the complete application of ultimate AGI in the real world. The main limitation of AGI in Level 1 and 2 is that the results of algorithms achieved on manual benchmarks and environments do not match the real world. The huge difference between the real environment and the benchmark is a huge challenge that affects the deployment of AGI in the real world. Super realistic simulation techniques make the deployment of AGI in the real world possible from the following aspects: 1) Realistic simulation can generate a large amount of high-quality data for AGI to perform self-training and self-evolving. Current benchmark data are often collected or designed by humans, and have noise and bias compared to real-world scenes. Realistic simulation based on some data-driven techniques like VAE (Kingma and Welling, 2013) and GAN (Goodfellow et al., 2020) can provide unbiased data to AGI to achieve better alignment. 2) AGI’s algorithms and strategies only need to be fine-tuned on the realistic simulator before they can be applied to the real world. Realistic simulators can not only simulate the interaction of different AGI agents in the real world but also reflect the causal laws of the real world. This allows the effects of AGI’s algorithms and strategies in the simulator to be replicated in the real world.
@@
  • A Call for rethinking and redefining the ultimate AGI. As we currently stand at the first stage of our AGI hierarchy, it is very possible that our understanding of higher-level AGI remains shaky or become outdated. Therefore, researchers might need to rethink and redefine what the ultimate AGI really is as we progress along the journey, the answer of which might depend on our gradually increased understanding of the difference between artificial and biological intelligence from both a biological and philosophical perspective, and could even be eventually limited by our current imagination. Once our understanding and goal change, a new set of evaluation frameworks and alignment procedures should be developed accordingly to meet the new expectations. It is worth keeping in mind the possibility that technical advancement might be “local” and people need to restore the wheels at some point in order to break the constraints towards AGI.
@@
The rapid development of AI has given people a lot of expectations and imaginations for more powerful AGI. In today’s era, analysis based on current AI development trends and deficiencies is an important way to measure our distance from AGI and how to achieve AGI.
@@
  • How far are we from AGI? Two valuable opinions are provided to discuss this topic. In Shane’s 2009 prediction, there is a 50% chance of AGl by 2028. In addition, Metaculus thinks that Al must pass a 2-hour, adversarial Turing test in text, images, and audio files, have robotic capabilities to assemble an automobile model, and high performance on difficult cognitive tests to achieve the first AGI system.
@@
From a probabilistic perspective, generalization depends largely on two properties of deep learning models, the support and the inductive biases. Starting from this, we can try to reason about whether we can build generally intelligent systems through the lens of Kolmogorov complexity and generalization bounds. Looking ahead, although there have been different signal showing the possibilities of building broadly intelligent systems, we might be still far away from doing that. In the future, we should embrace many safety considerations and alignment approaches when building these systems. Andrew Gordon Wilson introduces his views on how to build general intelligence as follows:
@@
  • Possibilities of building generalist models. Can we actually build “AGI” — which will be simultaneously good on many real-world problems? The no-free lunch theorems are sometimes used to argue that we can’t, which suggests we may need to build highly specialized learners for particular tasks. However, we think that universal learning (general intelligence) in the real world should be possible. Neural networks represent many compelling solutions to a given problem, which is perfect for Bayesian model averaging. Through the lens of Kolmogorov complexity, we can explore the alignment between structure in real-world data and machine learning models. A single low-complexity biased prior can suffice on a wide range of problems due to the low Kolmogorov complexity of data. Even under an arbitrarily large hypothesis space, generalization is possible if we assign prior mass disproportionately to the highly structured data that typically occurs. We can then design models that work well in small and large data regimes, by embracing a flexible hypothesis space combined with a strong simplicity bias.
@@
One of the fundamental questions that need to be addressed along the way toward AGI is how to relieve the tension between the demand and the supply of compute. One promise of AGI systems is to provide the service to everyone, which means that we need to serve the model on various devices, particularly on cheaper (e.g. lower memory capacity, worse compute capability) edge devices without sacrificing too much performance. Efficient AI computing, therefore, becomes one crucially important topic that tries to democratize AI for all users and devices. Song Han proposes two major versions of Edge AI as the step stones towards AGI as well as two ever-lasting questions that would help bring the realization of it:
  • Edge AI 1.0. The first category consists majorly of specialized models, usually trained with task-specific data, exhibiting limited generalization, and often still including failure of corner cases. Despite far from ideal, many works at this level have already shown impressive results in deploying models on resource-hungry platforms such as Efficient Inference Engine (EIE) Han et al. (2016) and Tiny ML that enables on-device pretraining of a model under 256KB memory which can still score decently on ImageNet (Lin et al., 2024f). This gives extremely promising direction for advancing towards the next stage.
@@
      Model compression to fight against the intrinsic limitation of limited memory on the device. Even with the existing of very strong base models like VILA, serving it on commodity or edge devices is non-trivial. LLM quantization stands out as a promising solution for this. SmoothQuant Xiao et al. (2024a), for example, smoothes out the activation outliers with a mathematically equivalent transformation for the ease of quantization based on empirical observation. AWQ (Lin et al., 2024b) quantize the LLM weights with activation awareness and a hardware friendly algorithm, which has been widely adopted by many popular frameworks to compress large models.
@@
      Efficient deployment to serve these quantized models both on the edge and over the cloud. TinyChat (tin, 2023; Lin et al., 2024b) provides both an efficient, light-weight, and python-naive framework to serve quantized LLMs and VLMs with low latency and great compatibility with other stacks. QServe (Lin et al., 2024c), on the other hand, targets the cloud deployment with quantization and system co-design, which can quantize the model up to 4-bit for efficient serving.
@@
      Long-context: StreamingLLM (Xiao et al., 2024b), along with the attention sink technique, enables long conversation within a non-stop streaming application, which primarily aims to reduce the extensive memory consumption and prevent perplexity explosion after exceeding the sequence length. Complimentary to this, LongLoRA (Chen et al., 2024b) solves the efficient long fine-tuning with specially designed shifted sparse attention pattern.

### v2

@@ 6.2 AGI Evaluation
The concept of evaluating AGI traces back to the famous Turing Test proposed by Alan Turing in 1950 (Turing, 1950). Turing posited that a machine could be considered intelligent if it could converse with a human so that the human could not distinguish whether they were conversing with a machine or another human. This laid the groundwork for the field of AGI evaluation. However, the Turing test has several drawbacks, such as its reliance on deception, subjective evaluation, and narrow focus on language use. To address these issues, a more comprehensive approach called the I-athlon (Adams et al., 2016) has been proposed, which evaluates machine intelligence across multiple dimensions and aims to provide a more objective and practical method for assessing progress in general-purpose AI.
@@ 6.2.1 Expectations for AGI Evaluation
It is undeniably challenging to design evaluation frameworks that satisfy all the recommended characteristics, but a well-constructed evaluation pipeline can help reflect the true power of increasingly sophisticated AGI systems.
@@
The new generation of model evaluations should focus on assessing AI systems across multiple dimensions, considering not only their ability to engage in human-like conversation but also their capacity to reason, learn, adapt, and solve complex problems. These tests should encompass a broader range of cognitive abilities (Lebiere, 2007) and evaluate the AI’s performance in real-world scenarios. Drawing from the concepts of AGI internal, external, and system levels as we discussed in earlier sections, we can outline the key aspects that a new AGI evaluation framework should address:
@@
  • System level. Examine the AI’s overall behavior and its ability to pursue long-term goals, collaborate with humans and other AI systems (Wang, 1006), and make ethical decisions. This could involve scenarios that assess the AI’s alignment with human values, its transparency and explainability, and its robustness and reliability in uncertain and adversarial situations.
@@
As we will discuss in the next section, the current evaluation frameworks are far away from achieving what we expect (Team, 2023). Here we list a couple of challenges associated with AGI evaluation (Xu and Ren, 2022) design, and we provide some concrete examples in each category:
@@
  • Long feedback loops. The trend toward more general-purpose AI indicates that models would gradually become more in people’s lives, making them more interactive. Instead of single-task solving, AGI systems will get multi-step feedback, often extending a longer period, making evaluation more complex. One commonly seen case happens in search engine development where a mature system needs to improve the click-through rate and track down users’ following actions such as purchases, comments, and after-sale satisfaction. Another potential situation appears when a health AI robot monitors a patient’s biological status constantly. Both of these require evaluations that span an elongated period.
  • Complicated environment setup. Many tasks inevitably presuppose more complicated environment setups such as robotic manipulation, self-driving assistance, and program synthesis. Scaling the evaluation of these applications necessitates a higher level of automatic environment generation, verification, and measurement. Often, the challenge associated with these tasks also comes with the difficulty of applying a single metric for comparison and can usually encounter many physical constraints (Peng et al., 2024b) that are hard to overcome.
  • Super-evaluation. Similar to super alignment, when AI models start to surpass humans, the evaluation becomes prohibitively more challenging. Taking the example of theorem proving, one can imagine that the real meaning of these models comes when they can prove unsolved theorems, which require greater expertise to verify. With the current setup, the framework might determine whether an AI agent can beat humans for a specific task (e.g., AlphaGo Silver et al. (2017) and  (Lake and Baroni, 2023)) but might not be as confident to access its absolute ability beyond. Fortunately, formal proof systems such as Lean (De Moura et al., 2015) may be helpful. Lean is an interactive proof system that utilizes formal logic to verify the correctness of mathematical theorems and computational outputs. As AI models begin to generate results that exceed human verification capabilities, systems like Lean become indispensable for ensuring the validity of these outputs.
@@ 6.2.2 Current Evaluations and Their Limitations
In this section, we summarize several representative works focusing on existing AI evaluation benchmarks. One category aims to provide a single pipeline suite consisting of many different tests such as OpenCompass (Contributors, 2023), AGIEVAL (Zhong et al., 2023a) and Huggingface Open LLM Leaderboard  [^9] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard for language understanding, reasoning, knowledge, and interaction abilities, and MT-BENCH (Zheng et al., 2024) for multi-task generalization ability. The GAIA benchmark (Mialon et al., 2023) constitutes a significant stride in this direction. It is specifically designed to evaluate General AI Assistants, presenting a series of real-world questions that test fundamental competencies such as reasoning, multi-modality handling, web browsing, and tool-use proficiency. AGENTBENCH is another comprehensive benchmark (Liu et al., 2024c) suite designed for evaluating the efficacy of LLMs as autonomous decision-making agents across eight interactive environments, highlighting the performance discrepancy between leading commercial models and open-source counterparts. On a granular level, there are many prior efforts in accessing a model’s specialized ability, which can be roughly divided into several aspects: accuracy, calibration and uncertainty, robustness, fairness, bias and stereotypes, toxicity, and efficiency (Liang et al., 2023), which can be further classified into two sets based on their objectives. OpenAGI (Ge et al., 2023b) is an open-source platform designed to advance AGI by integrating LLMs with domain-specific expert models. It utilizes a dual strategy of benchmark and open-ended tasks to evaluate.
@@
  • Creativity. Ever since the born of generative models, research has been ongoing into the use of them to model human creative processes, to mimic or complement them, in art, music, and literature (Cardoso et al., 2009; Colton et al., 2012). Creativity is mainly linked to the diversity of generated content with specific tasks like storytelling (Alhussain and Azmi, 2021) emerges. Recent works focus on understanding and prompting LLMs to generate creative textual content (Zhao et al., 2024c; Shanahan and Clarke, 2023).
@@
  • Lack of failure analysis. Almost all benchmarks we have seen so far give aggregated results, usually in the form of averages and percentiles. However, benchmarks should in principle provide more insights into improving a certain system. The most informative feedback would contain information about the analysis of the failing or worst cases. This can also showcase the potential pitfalls and risks associated with a specific system to help with the decision for danger-critical applications.
@@ 6.3 How to Get to the Next AGI Level
  • Automated Coding AI: Bridging the Gap to the Ultimate Artificial Intelligence Coding AI refers to AI systems capable of automatically planning and generating code to solve complex tasks. We believe that the advancement of such systems could significantly accelerate progress toward the Ultimate AGI. The AGI in levels 1 and 2 is limited since they require a variety of data collected by humans and require specific optimization goals designed by humans when they tackle different tasks. Human-in-the-loop makes it impossible for AGI at these two levels to realize self-evolve. Advanced coding AI solves the above challenges in two ways: 1) They enable AGI to interact with the real world and obtain large amounts of domain data. In the scenario of a single agent, an advanced coding AI takes writing code as the most basic tool for AGI to interact with the world, enabling AGI to plan and reason in the form of code and get feedback on the real world through the results of code. When it comes to multi-agent scenarios, each agent can be regarded as a unique coding AI based on their profile. Then, through the interactions between agents and the interaction between agents and the real world, AGI can obtain enough data for self-training and evolution. 2) It makes it possible for AGI to automatically define optimization goals for different tasks. With the ability to write codes, AGI can do try-and-error in the different tasks and obtain feedback through the interaction between code and the real-world environment. This feedback contains information about how well the AGI has completed its tasks and can take many forms, which can be differentiable or non-differentiable. Some techniques based on reinforcement learning can be introduced to use these different forms of feedback to align and self-evolve the AGI. In this case, the evolution of stronger AGI can be achieved without requiring humans to specify goals. More information about coding AI could be found on Sec 7.5
@@
  • Super realistic simulation promotes the complete application of ultimate AGI in the real world. The main limitation of AGI in Levels 1 and 2 is that the results of algorithms achieved on manual benchmarks and environments do not match the real world. The huge difference between the real environment and the benchmark is a huge challenge that affects the deployment of AGI in the real world. Super realistic simulation techniques make the deployment of AGI in the real world possible from the following aspects: 1) Realistic simulation can generate a large amount of high-quality data for AGI to perform self-training and self-evolving. Current benchmark data are often collected or designed by humans and have noise and bias compared to real-world scenes. Realistic simulation based on some data-driven techniques like VAE (Kingma and Welling, 2013) and GAN (Goodfellow et al., 2020), Transformers (Micheli et al., 2023), and Diffusion Models (Ding et al., 2024a; Alonso et al., 2024) can provide unbiased data to AGI to achieve better alignment. 2) AGI’s algorithms and strategies only need to be fine-tuned on the realistic simulator before they can be applied to the real world. Realistic simulators can not only simulate the interaction of different AGI agents in the real world but also reflect the causal laws of the real world. This allows the effects of AGI’s algorithms and strategies in the simulator to be replicated in the real world.
@@
  • A Call for rethinking and redefining the ultimate AGI. As we currently stand at the first stage of our AGI hierarchy, it is very possible that our understanding of higher-level AGI remains shaky or becomes outdated. Therefore, researchers might need to rethink and redefine what the ultimate AGI really is as we progress along the journey, the answer of which might depend on our gradually increased understanding of the difference between artificial and biological intelligence from both a biological and philosophical perspective, and could even be eventually limited by our current imagination. Once our understanding and goal change, a new set of evaluation frameworks and alignment procedures should be developed accordingly to meet the new expectations. It is worth keeping in mind the possibility that technical advancement might be “local” and people need to restore the wheels at some point in order to break the constraints towards AGI.
@@
The rapid development of AI has given people a lot of expectations and imaginations for a more powerful AGI. In today’s era, analysis based on current AI development trends and deficiencies is an important way to measure our distance from AGI and how to achieve AGI.
@@
  • How far are we from AGI? Two valuable opinions are provided to discuss this topic. In Shane’s 2009 prediction, there is a 50% chance of AGl by 2028. In addition, Metaculus thinks that Al must pass a 2-hour, adversarial Turing test in text, images, and audio files, have robotic capabilities to assemble an automobile model, and have high performance on difficult cognitive tests to achieve the first AGI system.
@@
From a probabilistic perspective, generalization depends largely on two properties of deep learning models, the support and the inductive biases. Starting from this, we can try to reason about whether we can build generally intelligent systems through the lens of Kolmogorov complexity and generalization bounds. Looking ahead, although there have been different signals showing the possibilities of building broadly intelligent systems, we might be still far away from doing that. In the future, we should embrace many safety considerations and alignment approaches when building these systems. Andrew Gordon Wilson introduces his views on how to build general intelligence as follows:
@@
  • Possibilities of building generalist models. Can we actually build “AGI” — which will be simultaneously good on many real-world problems? The no-free lunch theorems are sometimes used to argue that we can’t, which suggests we may need to build highly specialized learners for particular tasks. However, we think that universal learning (general intelligence) in the real world should be possible. Neural networks represent many compelling solutions to a given problem, which is perfect for Bayesian model averaging. Through the lens of Kolmogorov complexity, we can explore the alignment between structure in real-world data and machine learning models. A single low-complexity bias can suffice on a wide range of problems due to the low Kolmogorov complexity of data. Even under an arbitrarily large hypothesis space, generalization is possible if we assign prior mass disproportionately to the highly structured data that typically occurs. We can then design models that work well in small and large data regimes, by embracing a flexible hypothesis space combined with a strong simplicity bias.
@@
One of the fundamental questions that need to be addressed along the way toward AGI is how to relieve the tension between the demand and the supply of computing. One promise of AGI systems is to provide the service to everyone, which means that we need to serve the model on various devices, particularly on cheaper (e.g. lower memory capacity, worse compute capability) edge devices without sacrificing too much performance. Efficient AI computing, therefore, becomes one crucially important topic that tries to democratize AI for all users and devices. Song Han proposes two major versions of Edge AI as the step stones towards AGI as well as two ever-lasting questions that would help bring the realization of it:
  • Edge AI 1.0. The first category consists majorly of specialized models, usually trained with task-specific data, exhibiting limited generalization, and often still including failure of corner cases. Despite far from ideal, many works at this level have already shown impressive results in deploying models on resource-hungry platforms such as Efficient Inference Engine (EIE) Han et al. (2016) and Tiny ML that enables on-device pretraining of a model under 256KB memory which can still score decently on ImageNet (Lin et al., 2024f). This gives an extremely promising direction for advancing towards the next stage.
@@
      Model compression to fight against the intrinsic limitation of limited memory on the device. Even with the existence of very strong base models like VILA, serving it on commodity or edge devices is non-trivial. LLM quantization stands out as a promising solution for this. SmoothQuant Xiao et al. (2024a), for example, smoothes out the activation outliers with a mathematically equivalent transformation for ease of quantization based on empirical observation. AWQ (Lin et al., 2024b) quantizes the LLM weights with activation awareness and a hardware-friendly algorithm, which has been widely adopted by many popular frameworks to compress large models.
@@
      Efficient deployment to serve these quantized models both on the edge and over the cloud. TinyChat (tin, 2023; Lin et al., 2024b) provides both an efficient, lightweight, and python-naive framework to serve quantized LLMs and VLMs with low latency and great compatibility with other stacks. QServe (Lin et al., 2024c), on the other hand, targets the cloud deployment with quantization and system co-design, which can quantize the model up to 4-bit for efficient serving.
@@
      Long-context: StreamingLLM (Xiao et al., 2024b), along with the attention sink technique, enables long conversation within a non-stop streaming application, which primarily aims to reduce the extensive memory consumption and prevents perplexity explosion after exceeding the sequence length. Complimentary to this, LongLoRA (Chen et al., 2024c) solves the efficient long fine-tuning with specially designed shifted sparse attention pattern.